# Exact Maximin Share Fairness via Adjusted Supply

### v1

@@ 4 Fair Division of Chores
We will additionally consider instances wherein all the agents’ costs are additive and they rank the chores in the same order. That is, there exists an indexing of the $m$ chores $\{a_{1},a_{2},\ldots,a_{m}\}$ with the property that $c_{i}(a_{1})\geq c_{i}(a_{2})\geq\ldots\geq c_{i}(a_{m})$ for all the agents $i\in[n]$. We will refer to such settings as instances with additive ordered costs.
@@ 4.1 Monotone Costs
4.2 Additive Costs
@@
Proof.
$c_{i}(A^{\prime}_{i})\leq\frac{11}{9}\mu_{i}$ (17)
for each agent $i\in[n]$.
[…]
The theorem stands proved. ∎
4.3 Additive Ordered Costs
In this section we assume that cost functions are additive and that there exists an indexing $a_{1},a_{2},\dots a_{m}$ of the chores such that $c_{i}(a_{1})\geq c_{i}(a_{2})\geq\dots\geq c_{i}(a_{m})$ for all agents $i$. We show that for any fair division instance with such valuations, there exists a multi-allocation $\mathcal{A}$ such that the number of unassigned chores in $\mathcal{A}$ is $\widetilde{O}\left(\frac{m}{n^{1/4}}\right)$.
Theorem 4.4.
Every fair division instance $\langle[n],[m],\{c_{i}\}_{i=1}^{n}\rangle$ with chores and additive ordered costs admits an MMS multi-allocation $\mathcal{A}=(A_{1},\ldots A_{n})$ wherein the number of unassigned chores is at most $\frac{3m\left(\log m\right)^{3/2}}{n^{1/4}}$, i.e.,
@@ 4.3 Additive Ordered Costs
Algorithm 2 Pre-processing for additive chores
@@
Since the underlying instance $\mathcal{I}=\langle[n],[m],\{c_{i}\}_{i=1}^{n}\rangle$ consists of additive ordered costs, we can index the chores even in the reduced instance $\mathcal{I}^{\prime}=\langle N,U,\{c_{i}\}_{i\in N}\rangle$ such that the indexing conforms to the ordering induced by the costs uniformly for all the agents. Specifically, we will assume, without loss of generality, that the set $U$ consists of chores $a_{1},a_{2},\ldots,a_{|U|}$ indexed such that $c_{i}(a_{1})\geq c_{i}(a_{2})\geq\ldots\geq c_{i}(a_{|U|})$ for all the agents $i\in N$. We now reason about the number of chores assigned, with copies, under the random multi-allocation $\mathcal{R}$ and in any given sequence of chores $S=\{a_{r},a_{r+1},\ldots,a_{s-1},a_{s}\}$. We denote by $\mathcal{S}$ the collection of all chore sequences in $U$, i.e.,
@@
$|S|\ \left(1-\frac{\log m}{\alpha}\right)-\delta$ $=|S|-|S|\frac{\log m}{\alpha}-2\alpha\frac{|U|}{\sqrt{|N|}}\ \sqrt{\log m}$
$=S-\Delta.$
@@
2:  while there exist chores $a$ and $b$ with the properties that $a<b$ (i.e., $c_{i}(a)\geq c_{i}(b)$), $\chi^{\mathcal{A}^{\prime}}_{a}\geq 2$, and $\chi^{\mathcal{A}^{\prime}}_{b}=0$ do
3:     Let $h$ be the lowest-indexed chore with $\chi^{\mathcal{A}^{\prime}}_{h}\geq 2$ and let $\ell$ be the lowest-indexed chore such that $\ell>h$ and $\chi^{\mathcal{A}^{\prime}}_{\ell}=0$.
4:     Select an agent $i$ with $h\in A^{\prime}_{i}$ and update $A^{\prime}_{i}\leftarrow(A^{\prime}_{i}\setminus\{h\})\cup\{\ell\}$.
@@
Algorithm 3 Shifting Algorithm for Chores
@@
We will show that—given a multi-allocation $\mathcal{Q}$ that satisfies the stated guarantees with respect to chore sequences—Algorithm 3 returns the desired multi-allocation $\mathcal{A}^{\prime}$. Algorithm 3, in each iteration of its while-loop, replaces a chore $h$ in the selected agent $i$’s bundle, $A^{\prime}_{i}$, by another chore $\ell$ of lower cost. Since $h$ was previously allocated to at least two agents and $\ell$ was unallocated, the number of unallocated chores in the maintained multi-allocation $\mathcal{A}^{\prime}$ decreases by one in each iteration. That is, the algorithm terminates in polynomial time. Moreover, given that $h$ has a higher cost than $\ell$, the agents’ allocated costs $c_{i}(A^{\prime}_{i})$ are non-increasing throughout the execution of the algorithm. Hence, for the returned multi-allocation $\mathcal{A}^{\prime}$, in particular, we have $c_{i}(A^{\prime}_{i})\leq c_{i}(Q_{i})$, for all the agents $i$. This observation establishes property (i) as stated in the lemma.
@@
We next establish property (ii) by proving that upon termination of the algorithm, the returned multi-allocation $\mathcal{A}^{\prime}$ satisfies $\left\|\chi^{\mathcal{A}^{\prime}}\right\|_{z}\leq\Delta$. Assume, towards a contradiction, that $\left\|\chi^{\mathcal{A}^{\prime}}\right\|_{z}\geq\Delta+1$. Write $\ell_{0}$ to denote the chore of highest index with the property $\chi^{\mathcal{A}^{\prime}}_{\ell_{0}}=0$. Also, let $S=\{a\in U\mid a\leq\ell_{0}\}$ be the sequence of chores with index at most $\ell_{0}$. Since $\ell_{0}$ is the highest index chore that is not assigned ($\chi^{\mathcal{A}^{\prime}}_{\ell_{0}}=0$), we have $\chi^{\mathcal{A}^{\prime}}_{b}\geq 1$ for all chores $b$ with index higher than $\ell_{0}$. Hence, all the unassigned chores are in fact contained in $S$; recall that there are at least $(\Delta+1)$ unassigned chores under $\mathcal{A}^{\prime}$. Further, the fact that the while-loop of the algorithm terminated with multi-allocation $\mathcal{A}^{\prime}$ and $\chi^{\mathcal{A}^{\prime}}_{\ell_{0}}=0$ ensures that $\chi^{\mathcal{A}^{\prime}}_{a}\leq 1$ for all chores $a\in S$. These observations imply that
$\sum_{a\in S}\chi^{\mathcal{A}^{\prime}}_{a}\leq|S|-\Delta-1$ (24)
@@
Further, note that, since $\chi^{\mathcal{A}^{\prime}}_{\ell_{0}}=0$, during any iteration of the while-loop, the algorithm would not have replaced a copy of a chore $h\in S$ with a copy of a chore $\ell\notin S$. That is, it could not have been the case that, during any iteration, the considered chores $h$ and $\ell$ satisfy $h<\ell_{0}<\ell$, since such a relation between the indices contradicts the selection criterion of $\ell$ as the lowest-index chore with $\chi^{\mathcal{A}^{\prime}}_{\ell}=0$. Therefore, throughout the algorithm’s execution, for any considered chore $h\in S$, the corresponding chore $\ell$ must have also been contained in $S$. Considering this property along with the decrements and increments of $\chi^{\mathcal{A}^{\prime}}_{h}$ and $\chi^{\mathcal{A}^{\prime}}_{\ell}$, respectively, we obtain $\sum_{a\in S}\chi^{\mathcal{Q}}_{a}=\sum_{a\in S}\chi^{\mathcal{A}^{\prime}}_{a}$.
@@
The last equality and equation (24) give us $\sum_{a\in S}\chi^{\mathcal{Q}}_{a}\leq|S|-\Delta-1$. This bound, however, contradicts the condition provided for $\mathcal{Q}$ in the lemma statement. Hence, by way of contradiction, we obtain property (ii), $\left\|\chi^{\mathcal{A}^{\prime}}\right\|_{z}\leq\Delta$. The lemma stands proved. ∎
@@ 4.4 Lower Bound for Chores
We now show that there exists fair division instances, with $m$ chores and monotone costs, such that in every MMS multi-allocation, at least $(1-o(1))\nicefrac{{m}}{{e}}$ chores remain unassigned. This lower bound shows that the positive result obtained in Theorem 4.2 is essentially tight.
@@
5 Conclusion
In this work, we considered a relaxation of the fair item allocation problem wherein it is permissible to make post facto adjustments to the supply of indivisible items towards achieving exact MMS fairness. In particular, we showed that limited duplication of goods or limited disposal of chores renders MMS fairness feasible, even for instances with monotone valuations/costs. We also demonstrated improved bounds for the special cases of additive and additive ordered valuations. It is worth remarking that our model frames the distinction between fair allocation of goods and that of chores in terms of the distinction between vector packing and covering problems. Indeed, finding multi-allocations with limited duplication of goods corresponds to packing copies of goods into agents’ bundles while satisfying some capacity constraints. Similarly, finding multi-allocations with limited disposal of chores corresponds to distributing chores among agents while covering as many chores as possible.
An interesting question is to study the strategic behaviour of agents in our model – is it possible to design a truthful mechanism that accepts as input each agent’s MMS-inducing partitions and outputs an MMS multi-allocation with bounded supply adjustment? Another direction for future work is to consider our framework of supply adjustments for other fairness notions that are either infeasible or not known to be feasible. For envy-freeness up to any good (EFx), such an approach would complement prior work on EFx with charity [9], which proves existential guarantees for EFx with disposal, rather than duplication, of goods.

### v2

@@ 3.4 Lower Bound for Goods
Remark.
We note that, for exact MMS, the $\ell_{\infty}$ lower bound given in Theorem 3.10 holds even under XOS valuations. In particular, as in the proof of Theorem 3.10, we select independently for each agent $i\in[n]$, a partition $\mathcal{P}^{i}=(P^{i}_{1},P^{i}_{2},\dots,P^{i}_{n})$ uniformly at random from the set $\Pi_{n}([m])$ of all possible $n$-partitions of $[m]$. Then, for each agent $i\in[n]$ and subset $S\subseteq[m]$, we set the valuation
$v_{i}(S)\coloneqq\max_{1\leq j\leq n}\ \frac{1}{|P^{i}_{j}|}|S\cap P^{i}_{j}|.$
Note that the valuations $v_{i}$ are point-wise maximizers of additive functions of the form $S\mapsto\frac{1}{|P^{i}_{j}|}|S\cap P^{i}_{j}|$. Hence, $v_{i}$s are XOS. Further, $v_{i}(S)=1$ if $P^{i}_{j}\subseteq S$, for some index $j\in[n]$, and $v_{i}(S)<1$ otherwise. Therefore, the partition $\mathcal{P}^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ certifies that, for each agent $i\in[n]$, the maximin share $\mu_{i}=1$. Moreover, for any MMS multi-allocation $\mathcal{A}=(A_{i},\dots,A_{n})$ it must hold that $P^{i}_{j}\subseteq A_{i}$ for each agent $i\in[n]$ and some index $j\in[n]$. This is exactly the property we utilize for the proof of Theorem 3.10, and the rest of the proof relies solely on this condition. Therefore, the lower bound holds in particular for XOS valuations.
@@ 4 Fair Division of Chores
We will additionally consider (in Section 4.2) instances with cost functions that are identically ordered.
@@ 4.1 Monotone Costs
4.2 Identically Ordered Costs
This section establishes upper bounds on supply adjustments under cost functions are identically ordered; the setup here is analogous to the one given in Section 3.2 for goods. Formally,
Definition 4.3.
In a fair division instance $\langle[n],[m],\{c_{i}\}_{i=1}^{n}\rangle$, the cost functions are said to be identically ordered if there exists an indexing $a_{1},a_{2},\ldots,a_{m}$ over the chores such that, for any pair of chores $a_{s},a_{t}$, with index $s<t$, we have $c_{i}(S+a_{s})\geq c_{i}(S+a_{t})$ for all agents $i\in[n]$ and each subset $S$ that does not contain $a_{s}$ and $a_{t}$.
We show that for any fair division instance with identically ordered costs, there exists a multi-allocation $\mathcal{A}$ such that the number of unassigned chores in $\mathcal{A}$ is $\widetilde{O}\left(\frac{m}{n^{1/4}}\right)$.
@@ 4.2 Identically Ordered Costs
Every fair division instance $\langle[n],[m],\{c_{i}\}_{i=1}^{n}\rangle$ with chores and identically ordered costs admits an MMS multi-allocation $\mathcal{A}=(A_{1},\ldots A_{n})$ wherein the number of unassigned chores is at most $\frac{3m\left(\log m\right)^{3/2}}{n^{1/4}}$, i.e.,
@@
Algorithm 2 Pre-processing for Identically Ordered Costs
@@
Since the underlying instance $\mathcal{I}=\langle[n],[m],\{c_{i}\}_{i=1}^{n}\rangle$ consists of identically ordered costs, this property continues to hold for the reduced instance $\mathcal{I}^{\prime}=\langle N,U,\{c_{i}\}_{i\in N}\rangle$ as well. Specifically, we will assume, without loss of generality, that the set $U$ consists of chores $a_{1},a_{2},\ldots,a_{|U|}$, indexed such that for any $a_{s}$ and $a_{t}$, with $s<t$, we have $c_{i}(S+a_{s})\geq c_{i}(S+a_{t})$, for all agents $i\in N$ and all subsets $S\subset U$ that do not contain $a_{s}$ and $a_{t}$. We now reason about the number of chores assigned, with copies, under the random multi-allocation $\mathcal{R}$ and in any given sequence of chores $S=\{a_{r},a_{r+1},\ldots,a_{s-1},a_{s}\}$. We denote by $\mathcal{S}$ the collection of all chore sequences in $U$, i.e.,
@@
$|S|\ \left(1-\frac{\log m}{\alpha}\right)-\delta$ $=|S|-|S|\frac{\log m}{\alpha}-2\alpha\frac{|U|}{\sqrt{|N|}}\ \sqrt{\log m}=S-\Delta.$
@@
2:  while there exist chores $b$ and $b^{\prime}$ with the following properties: (i) index of $b$ is less than that of $b^{\prime}$ (i.e., $b$ has higher marginal costs than $b^{\prime}$), (ii) $\chi^{\mathcal{A}^{\prime}}_{b}\geq 2$, and (iii) $\chi^{\mathcal{A}^{\prime}}_{b^{\prime}}=0$ do
3:     Let $a_{s}$ be the lowest-indexed chore with $\chi^{\mathcal{A}^{\prime}}_{a_{s}}\geq 2$ and let $a_{t}$ be the lowest-indexed chore such that $t>s$ and $\chi^{\mathcal{A}^{\prime}}_{a_{t}}=0$.
4:     Select an agent $i$ with $a_{s}\in A^{\prime}_{i}$ and update $A^{\prime}_{i}\leftarrow(A^{\prime}_{i}\setminus\{a_{s}\})\cup\{a_{t}\}$.
@@
Algorithm 3 Copy Redistribution for Identically Ordered Costs
@@
We will show that—given a multi-allocation $\mathcal{Q}$ that satisfies the stated guarantees with respect to chore sequences—Algorithm 3 returns the desired multi-allocation $\mathcal{A}^{\prime}$. Algorithm 3, in each iteration of its while-loop, replaces a chore $a_{s}$ in the selected agent $i$’s bundle, $A^{\prime}_{i}$, by another chore $a_{t}$ of higher index $t>s$. Since $a_{s}$ was previously allocated to at least two agents and $a_{t}$ was unallocated, the number of unallocated chores in the maintained multi-allocation $\mathcal{A}^{\prime}$ decreases by one in each iteration. That is, the algorithm terminates in polynomial time. Moreover, given that $a_{t}$ has a higher marginal cost than $a_{s}$ with respect to $A^{\prime}_{i}\setminus\{a_{s}\}$, the agents’ allocated costs $c_{i}(A^{\prime}_{i})$ are non-increasing throughout the execution of the algorithm. Hence, for the returned multi-allocation $\mathcal{A}^{\prime}$, in particular, we have $c_{i}(A^{\prime}_{i})\leq c_{i}(Q_{i})$, for all the agents $i$. This observation establishes property (i) as stated in the lemma.
@@
We next establish property (ii) by proving that upon termination of the algorithm, the returned multi-allocation $\mathcal{A}^{\prime}$ satisfies $\left\|\chi^{\mathcal{A}^{\prime}}\right\|_{z}\leq\Delta$. Assume, towards a contradiction, that $\left\|\chi^{\mathcal{A}^{\prime}}\right\|_{z}\geq\Delta+1$. Write $a_{r}$ to denote the chore of highest index $r$ with the property $\chi^{\mathcal{A}^{\prime}}_{a_{r}}=0$. Also, let $S$ denote the sequence of chores in $U$ with index at most $r$. Since $a_{r}$ is the highest index chore that is not assigned ($\chi^{\mathcal{A}^{\prime}}_{a_{r}}=0$), we have $\chi^{\mathcal{A}^{\prime}}_{b^{\prime}}\geq 1$ for all chores $b^{\prime}$ with index higher than $r$. Hence, all the unassigned chores are in fact contained in $S$; recall that there are at least $(\Delta+1)$ unassigned chores under $\mathcal{A}^{\prime}$. Further, the fact that the while-loop of the algorithm terminated with multi-allocation $\mathcal{A}^{\prime}$ and $\chi^{\mathcal{A}^{\prime}}_{a_{r}}=0$ ensures that $\chi^{\mathcal{A}^{\prime}}_{b}\leq 1$ for all chores $b\in S$. These observations imply that
$\sum_{b\in S}\chi^{\mathcal{A}^{\prime}}_{b}\leq|S|-\Delta-1$ (22)
@@
Further, note that, since $\chi^{\mathcal{A}^{\prime}}_{a_{r}}=0$, during any iteration of the while-loop, the algorithm would not have replaced a copy of a chore $a_{s}\in S$ with a copy of a chore $a_{t}\notin S$. That is, it could not have been the case that, during any iteration, the considered chores $a_{s}$ and $a_{t}$ satisfy $s<r<t$, since such a relation between the indices contradicts the selection criterion of $a_{t}$ as the lowest-index chore with $\chi^{\mathcal{A}^{\prime}}_{a_{t}}=0$. Therefore, throughout the algorithm’s execution, for any considered chore $a_{s}\in S$, the corresponding chore $a_{t}$ must have also been contained in $S$. Considering this property along with the decrements and increments of $\chi^{\mathcal{A}^{\prime}}_{a_{s}}$ and $\chi^{\mathcal{A}^{\prime}}_{a_{t}}$, respectively, we obtain $\sum_{b\in S}\chi^{\mathcal{Q}}_{b}=\sum_{b\in S}\chi^{\mathcal{A}^{\prime}}_{b}$.
@@
The last equality and equation (22) give us $\sum_{b\in S}\chi^{\mathcal{Q}}_{b}\leq|S|-\Delta-1$. This bound, however, contradicts the condition provided for $\mathcal{Q}$ in the lemma statement. Hence, by way of contradiction, we obtain property (ii), $\left\|\chi^{\mathcal{A}^{\prime}}\right\|_{z}\leq\Delta$. The lemma stands proved. ∎
@@ 4.3 Lower Bound for Chores
We now prove that there exists fair division instances, with $m$ chores and monotone costs, such that in every MMS multi-allocation, at least $(1-o(1))\nicefrac{{m}}{{e}}$ chores remain unassigned. This lower bound shows that the positive result obtained in Theorem 4.2 is essentially tight.
@@
4.4 Additive Costs
Theorem 4.9.
Proof.
$c_{i}(A^{\prime}_{i})\leq\frac{11}{9}\mu_{i}$ (26)
for each agent $i\in[n]$.
[…]
The theorem stands proved. ∎
5 Conclusion and Future Work
This work establishes that exact MMS fairness can be guaranteed in settings that permit post facto adjustments to the supply of indivisible items. We obtain tight bounds for general monotone valuations and monotone costs. Our specialized bounds for identically ordered valuations and costs bring forth novel applications of concentration inequalities in the fair division context. The techniques also highlight a distinction between goods and chores in the MMS context – for the goods case, the problem in the current context corresponds to a vector packing problem. By contrast, in the chores case we have a vector covering problem.
The framework of adjusted supply provides a meaningful route for bypassing barriers in discrete fair division. In particular, it would be interesting to develop truthful mechanisms for MMS that leverage bounded supply adjustment. Another relevant direction for future work is to consider the framework for achieving envy-freeness up to any good (EFX). Such an approach would complement prior works on EFX with charity (see, e.g., [9]), which obtain existential guarantees for EFX with disposal, rather than duplication, of goods.